Your Corporation has the following number of shares of stock outstanding: Cumulative Preferred stock, 2.5%. $100 par 8,000 shares Common stock,$10 par 10,000 shares It will distribute $100,000 to the two classes of stockholders this year. Including the current year, the preferred stock is three years in arrears.

How much of the dividend will the common stockholders receive?

$100,000

$ 20,000

$ 40,000

$ 60,000

Your Corporation is authorized to issue 800,000 shares of $3.00 par value common stock. When the company decided to issue a stock dividend, the market price of the stock was $15 per share. Book value was $8. Prior to the stock dividend, the stockholders' equity accounts report the following balances:

Common stock, $3 par, 800,000 shares authorized, 153,000 shares issued and 145,000 outstanding

                                           $   459,000

Paid in capital in excess of par

     630,000

Treasury stock, 8,000 shares $ 10

(     80,000)

Retained earnings

     785,000

Total stockholders’ equity

$1,794,000

What is the dollar value of a 5% stock dividend?

$114,750

$21,750

$108,750

$22,950
Answer #1

1)Correct option is "C"

In case of Cumulative preferred stock,any dividend in arrears is paid in preference to current year preferred and current stock dividend

Preferred dividend per year =Par value * number of shares* preferred dividend rate

              = 100*8000 *2.5%

             = 20000 per year

Total preferred dividend (including current year ) for 3 years= 20000*3=60000

Dividend to current stockholders =Total dividend- preferred dividend

              = 100000-60000

             = 40000

2)correct option is "C"-108750

Number of shares issued as stock dividend =number of shares outstanding *stock dividend

                                = 145000 * 5%

                                = 7250 shares

value of stock dividend=Number of shares issued as stock dividend*market price

                   = 7250 *15

                  = $ 108750